OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

Directrix (directrix):

Let A = the number of adults going on the trip Let S = the number of students going on the trip Then, A + S = 150 Each adult ticket cost $45 so the total amount of money the adults spend on tickets is $45 times A. Each student ticket cost $28.50 so the total amount of money the students spend on tickets is $28.50 times S. Solve these two equations simultaneously: A + S = 150 45*A + 28.5*S = 4770 --------------------

OpenStudy (anonymous):

hope this helps :) http://openstudy.com/study#/updates/4fe0ade9e4b06e92b86fd201